For math class, the passing rate is 53%. Find the probability that at least 23 students pass your course.sample size 32 mean 64 15.82 Sample standard deviation population standard deviation 12.5 max 90 Q1 60 Median 69.5 Q3 75

Please write notes readable.
Answer #1

P (Parsing rate) = 53010 - 0,53 9 1-P = 1-0.53 = 0.47 na 32 its a Binominal probability formula : ncoope.q we want probabilit